OpenStudy (anonymous):

put 1/2y=x-1 into y=mx+b? please show how?

OpenStudy (anonymous):

y=2x-2

OpenStudy (aroub):

oh just put 1/2y=x-1 in the form y=mx+b like : y=mx+b 1/2y=x-1 (u have to get rid of the 2 , so u multiply both sides by 2) y=2x-2 did u understand??

OpenStudy (anonymous):

yes, I got it, thank you so much. :)

OpenStudy (aroub):

no problem!1 :)